OpenStudy (anonymous):

find the absolute minimum point for the graph of f(x)=x^4+3x^3+x^2-1

OpenStudy (anonymous):

This is a question testing on the chapter Differentiation let f(x) = y Y= x^4 + 3x^3 + x^2 -1 DY/DX = 4x^3 + 9x^2 + 2x For turning pt so called min/max pt, DY/DX = O 4x^3 + 9x^2 + 2x = o x(4x^2 + 9x + 2) = o Find the coordinates of x and sub it into the first y equation at the above. They are the coordinates of the min pt.

OpenStudy (anonymous):

i dont know how to do that cause when my teacher taught it i was out on maternity leave

OpenStudy (anonymous):

Now that you saw the solution to the question, you should memorise it as this is not a topic about understanding. Hope it helps you :)

OpenStudy (anonymous):

it doesnt but thanks for trying